Determine the partial differential equation arising from

ax2+by2+z2=4


1
Expert's answer
2021-11-02T11:59:01-0400

We can write the given relations as:


"f(x, y, z)=ax^2+by^2+z^2-4=0"

then differentiating partially with respect to "x" and "y" respectively, we have


"\dfrac{\partial f}{\partial x}+\dfrac{\partial f}{\partial z}\cdot\dfrac{\partial z}{\partial x}=0 \text{ (Keeping }\dfrac{\partial y}{\partial x}=0)"

"\dfrac{\partial f}{\partial y}+\dfrac{\partial f}{\partial z}\cdot\dfrac{\partial z}{\partial y}=0 \text{ (Keeping }\dfrac{\partial x}{\partial y}=0)"

Or


"2ax+2z\cdot\dfrac{\partial z}{\partial x}=0 =>ax+zp=0"

"2by+2z\cdot\dfrac{\partial z}{\partial y}=0 =>by+zq=0"

"a=-\dfrac{zp}{x}, b=-\dfrac{zq}{y}"

Substitute


"-\dfrac{zp}{x}x^2-\dfrac{zq}{y}y^2+z^2-4=0"

"-pxz-qyz+z^2-4=0"

which is required partial differential equation.